OpenStudy (hba):

2x+3y=6 can be expressed in terms of conjugate coordinates as :

OpenStudy (hba):

@thivitaa

OpenStudy (hba):

@campbell_st

OpenStudy (campbell_st):

well it doesn't make a lot of sense to me the slope intercept form is y = -2/3x + 2

OpenStudy (hba):

Let me give you the options @campbell_st z+3/2i(z-zbar)=6,zbar+3/2i(z-zbar)=6,z+zbar-3/2i(z-zbar)=6,z-zbar+3/2i(z-zbar)=6

OpenStudy (hba):

So @thivitaa

OpenStudy (thivitaa):

cant gt the ans....bad at math..:)

OpenStudy (hba):

Ok

OpenStudy (hba):

@dpaInc

OpenStudy (hba):

@sirm3d

OpenStudy (sirm3d):

what does conjugate coordinates mean?

OpenStudy (hba):

Have You ever studied Complex numbers ?

OpenStudy (sirm3d):

yes i did. and i can't remember conjugate coordinates. can you refresh me on this?

OpenStudy (hba):

Ok,Any pair of complex numbers a+ib and a-ib have a product which is real,since (a+ib)(a-ib)=a^2-abi+abi-b^2i^2=a^2+b^2. Such complex numbers are called "conjugate" and each is the conjugate of the other.If a+ib is denoted by z,then its conjugate a-ib,is denoted by zbar. I hope this helps.

OpenStudy (sirm3d):

@hba i get it. the answer is \[\huge z+z^b-\frac{ 3 }{ 2i }(z-z^b)=6 \] where z^b is the conjugate of z. for the details. if we let z=x+yi and z^b = x-yi then \[\huge x=\frac{ z+z^b }{ 2 }\] and \[\huge y=\frac{ z-z^b }{ 2i }\] just replace x and y the their equivalent expressions